Please help me with this problem. Thanks Sam & Tony purchase a house. They find one that is selling for $450,000. They have been advised to make a 20% down payment. How much total interest will they pay if they take the full 20 years to pay off the mortgage with an interest rate of 3.75%? Please show the formula/work, Thanks

R=Monthly Payment

P= Principal

r- rate

n- number of payments per year

t=number of years
Answer #1

selling price is $450000

a down payment is 20%

so loan amount is

L= 450000 – 0.2 * 450000

L = 450000 - 90000

L = 360000

.

Y = 1 (5+1)-1

L=

1-(1 + 0.0375)-12-20 360000 = R 0.0375

360000 = R 1-(1+0.003125) -240 0.003125

360000 = R 1- (1.003125) -240 0.003125

360000 = R 1-0.4729 0.003125

360000 = R 0.52708 0.003125

360000 = R186.6658

R=\frac{360000}{186.6658}

{\color{red} R= 2134.39793}..................Monthly Payment

.

.

.

total interest is given by

1 = Rn.t-1

I= 2134.39793\cdot 12\cdot 20-360000

I = 512255.5032 - 360000

{\color{Red} I=152255.5032}.................. total interest